OpenStudy (anonymous):

r^2+13r-30=0 step by step please thanks

OpenStudy (anonymous):

First step is to factor the polynomial.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

whats the polynomial?

OpenStudy (anonymous):

r^2+13r-30

OpenStudy (anonymous):

r^2+13r-30 is the polynomial. A term is something like r^2 or 13r. when you have more than one term or two terms, or in this case 3 terms, it's a polynomial. SO the whole thing is a polynomial.

OpenStudy (anonymous):

i could have sworn we just did this one...

OpenStudy (anonymous):

and the way to find how r is = 0 is to use the factor method. so set up two blank sets of ( ) ( ) = 0

OpenStudy (anonymous):

yeah but i asked questions no one responded i needed it explained further sorry

OpenStudy (anonymous):

we did satellite. It's he needs explained further :)

OpenStudy (anonymous):

oops ok sorry

OpenStudy (anonymous):

so when you have ( ) ( ) = 0, you take your polynomial and go hmm. With all polynomails begging with something like r^2, you are going to need two r's in the ( ) ( )

OpenStudy (anonymous):

Oh I remembered learning this lol just forgot how to, Im doing summer math problems and need my memory refreshed lol

OpenStudy (anonymous):

so it should at first look like this, (r ) ( r ) = 0 I'm doing summer math too.

OpenStudy (anonymous):

so those two R's are going to multiply in the end to make r^2, that's why this works. next you need two factors of 30 that multiply to make -30 and add to make 13r

OpenStudy (anonymous):

oh yeah. so it's supposed to be equaled to the coefficient?

OpenStudy (anonymous):

in the equation

OpenStudy (anonymous):

yup, so r x r = r^2. So all the factors of 30? 30 and 1, 15 and 2, 5 and 6 ect

OpenStudy (anonymous):

of those factors, can you see any two that would add or subtract to be 13?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

5 + 6 = 11. :P but 15-2 = 13

OpenStudy (anonymous):

oohhh lol oops sorry

OpenStudy (anonymous):

I remember now thanks so much for refreshing my memory xD

OpenStudy (anonymous):

np XD. so if you know that 15-2 = 13 and 15x2 = 30 then you know what two numbers go in the ( ) ( ). So you should have (r 15) and (r 2) now. All that's left is to find which sign (eithr + or -) to put between!

OpenStudy (anonymous):

no prob. take it easy. I really want subway right now. I should work on my precalc stuff

OpenStudy (anonymous):

lol okay thanks again

OpenStudy (anonymous):

im going to refer back to your explanation on the rest of these problems

OpenStudy (anonymous):

are you there?